OpenStudy (anonymous):

The slope of the line passing through P(1,4) and Q(x,1) is -3/5. What is the value of x?

OpenStudy (lukecrayonz):

y2-y1/x2-x1

OpenStudy (lukecrayonz):

1-4/x-1=-3/5

OpenStudy (lukecrayonz):

-3/x-1=-3/5

OpenStudy (anonymous):

that may be correct lls

OpenStudy (anonymous):

The answer choices are: A. 6 B. 7 C. 8 D. 9 E. 10

OpenStudy (anonymous):

the answer i got is 6.666667 so would it be A or B?

OpenStudy (anonymous):

Luke took you most of the way -- you don't even really need to do anything from here - you should be able to tell by looking at it: \[\frac{ -3 }{ x-1 }=\frac{ -3 }{5 }\]

OpenStudy (anonymous):

so you cross multiple

OpenStudy (anonymous):

You could -- but if you look, you'll see both numerators are the same, so you just need to set the denominators equal to each other. (x-1 = 5)

OpenStudy (anonymous):

Sure. Actually, I started out suggesting you cross-multiply, but when I was making the fraction in the equation editor, I noticed that I could solve it just by looking at it instead.

OpenStudy (anonymous):

ok. well thanks a lot